OpenStudy (anonymous):

Can anyone tell me if this is set up correct? find rate of change 3x^2+6 f(4)-f(2)/4-2 f(4)=4(4)^2+4+4+2=26 OR f(4)=f(4)^6+4+4+6=4110

OpenStudy (anonymous):

rate of change is from 2 to 4

OpenStudy (anonymous):

f(4) = 3(4)^2 + 6 = 3(16) + 6 = 54 f(2) = 3(2)^2 + 6 = 3(4) + 6 =18 then~ (f(4)-f(2))/(4-2)